OpenStudy (superhelp101):

How many grams of nitric acid will react completely with a block of iron metal that is 4.5 cm by 3.0 cm by 3.5 cm, if the density of iron is 7.87 g/mL? Show all steps of your calculation as well as the final answer. Fe + HNO3 → Fe(NO3)3 + H2

OpenStudy (superhelp101):

i know i have to do this first i have to balance the equation and then you can get mass as d=m/V

OpenStudy (abmon98):

Multiply 4.5, 3 and 3.5 to obtain the volume of the block of iron. Density=Mass/Volume Density*Volume=Mass

OpenStudy (superhelp101):

oh. after i do do Density*Volume=Mass do get the answer?

OpenStudy (abmon98):

yes you do get the mass of Fe

OpenStudy (superhelp101):

Volume: 4.5 X 3 X 3.5= 47.25 Density*Volume=Mass 7.87 *47.25= 371.8575 Density=Mass/Volume 7.87=371.8575/ 47.25 7.87=7.87 :)

OpenStudy (superhelp101):

how do i find how many grams?

OpenStudy (abmon98):

mass is known now, use Number of Moles=Mass(g)/Molar Mass(g/mol) plug in the value of 371.8575 as the mass and look for the atomic weight of iron on your periodic table and plug it in as the molar mass Balance your chemical equation, look for the mole ratio calculate the molar mass of nitric acid and use Mass(g)=Molar Mass(g/mol)*Number of Moles

OpenStudy (superhelp101):

ok i will do it and then i'll let you know what i get

OpenStudy (abmon98):

take your time.

OpenStudy (superhelp101):

Moles=Mass(g)/Molar Mass(g/mol) 6.66=371.8575/ 55.85 Balance equation is Fe2O3 + 6HNO3 → 2Fe(NO3)3 + 3H2O Mass(g)=Molar Mass(g/mol)*Number of Moles 371.8575= 55.85 * 6.66 371.8575= 371.961 it is about the same did i do anything wrong?

OpenStudy (abmon98):

now you got the number of moles of iron oxide look at the mole ratio of Fe2O3 to HNO3 which is 1:6 multiply out 6.66 by 6 and divide by 1

OpenStudy (superhelp101):

so 6.66 * 6=39.96/1= 39.96

OpenStudy (superhelp101):

@Abmon98

OpenStudy (abmon98):

no 39.96 represent the number of moles of HNO3 now multiply 39.96 by the molar mass of HNO3 1+14+(16*3)

OpenStudy (superhelp101):

39.96 * 63.01= 2517.8796

OpenStudy (superhelp101):

@Abmon98

OpenStudy (abmon98):

Exactly

OpenStudy (superhelp101):

yay so would this be the answer 2517.8796 for grams?

OpenStudy (abmon98):

yes

OpenStudy (superhelp101):

ok. thank you so much for your time and help!

OpenStudy (abmon98):

your most welcome man and good to see such improvement :D